### v1

# Zhentao Zhang

Abstract
It is well known that a constant energy density of the vacuum produces a negative pressure. We show that, in the presence of special boundaries for the photon field, the virtual photons in the quantum vacuum may induce a negative pressure of this type. In this context, we find in particular that an anomalous quantum vacuum radiation is responsible for the appearance of this negative pressure in a system, which generates the tangential Casimir force experienced by parallel conducting plates.
The cosmological constant in Einstein’s equations is equivalent to the existence of a constant vacuum energy density [1]. Inevitable quantum fluctuations imply that the vacuum energy may originate from the zero-point energy in the vacuum, and this perspective is encouraged by the overall positive sign for the reality of the quantum vacuum [2]. Aside from its profound cosmological implications, itself has a peculiar property that it produces a negative pressure. The construction of the energy-momentum tensor of the vacuum could provide an argument for this property [3]. And essentially, the negative pressure is an effect required by the conservation of energy. However, it is of interest to understand the appearance of negative pressure from substructure of the quantum vacuum; be that as it may, it is hard to imagine that such an understanding can be fully achieved at this stage, since we do not yet understand the huge discrepancy between the observational data of the dark energy [4] and the prediction of zero-point energy. Therefore, in this work we focus on a simpler but well-defined variant of this negative pressure problem: We shall show that an observable negative pressure in the vacuum could occur in confined spaces for the photon field and that its appearance is a result of the quantum vacuum radiation, which may offer some insights into the quantum origin of the negative pressure.
Let us consider the system in Fig. 1. In addition to the normal Casimir force, there is an attractive tangential force between the plates [5]
along the -direction. A thorough discussion on the presence of this force in the configuration was given in Ref. [5], which follows the conventional mechanism of the modification of zero-point energy. However, as briefly remarked by Casimir himself [6] and explored in detail by Ref. [7], there is a more direct physical interpretation of the Casimir effect: the normal Casimir force may be a consequence of the radiation pressure exerted by the virtual/zero-point photons in the quantum vacuum, see also an investigation in a different context [8]. To achieve the main aim of the present study, we shall show first that the radiation mechanism may also induce the tangential force in Eq. (1), where an anomaly will be detected.
Figure 1: A side view for the parallel, perfectly conducting semi-infinite plates separated by the vacuum and overlapped partially along the -direction in the Cartesian coordinate system. The overlap distance is much larger than the separation , and the infinite width of the plates along the -direction will be denoted by .
The radiation interpretation comes from the fact that each virtual photon in the quantum vacuum carries a zero-point momentum [9] and the incident field of a plane wave reflecting off a surface exerts a pressure [7]
on the surface, where is the wave vector, is the energy density of the incident field, and is the angle of incidence.
In the misaligned system, the configuration imposes a constraint at the overlapping edges for the incident field. Only those virtual photons that carry the momenta satisfying can pass into the space between the parallel plates. The other virtual photons reflect off the edge and then a radiation pressure occurs.
Assume is the zero-point energy density of a mode of frequency in a quantization volume . By
for each mode, removing the superfluous contribution from the modes with , the radiation pressure on the left boundary (say) is found to be
where
and
where the quantization volume for is , the summation includes a weight factor for the term, and a factor of 2 for the two independent polarizations is included.
After some straightforward algebra, the radiation pressure may be written in the form of
Introduce a standard Casimir-type cutoff function that would make each integrand in the integral decays sufficiently fast for large , and its derivatives
Then, applying the Euler-MacLaurin formula to Eq. (7), we may find
It is seen that the result is negative, although should be an effect induced by the reflection of the virtual photons. Hence, we observe an anomaly in the sense that the virtual photons reflecting off a surface exert effectively a negative pressure on the surface. Note that, to avoid ambiguity of the term of negative pressure which is associated with a constant energy density, we will call this negative radiation pressure the anomalous pressure.
The anomalous radiation pressure should affect only the upper plate in the system since its tangential displacement causes the displacement of the left boundary. Thus, the pressure on the cross section of generates a force
acting on the upper plate, which is exactly the attractive force in Eq. (1) [The minus in Eq. (12) indicates that the direction of the force is opposite to the -axis]. Clearly, the physics of the right boundary would be the same.
However, it is crucial to realize that the misaligned system may be an unusual physical realization, in view of the fact that the tangential motion of the plates is precisely equivalent to a physical process that the expansion or contraction of a space with a finite, positive constant vacuum energy density . Hence, the system implies a negative pressure on the side boundary as
where is the vacuum energy of a volume with the finite constant energy density, and the change of the volume is determined by the change of in the misaligned system. The pressure is identical to Eq. (11). Therefore, we may conclude that the anomalous quantum vacuum radiation could be responsible for the appearance of the negative pressure in the configuration.
The above discussions might lead one to suppose that the anomalous radiation may be a universal mechanism to generate the negative pressure in the quantum vacuum. However, as we shall show below, this may not be true.
Consider that in the misaligned system one conducting plate is replaced with an infinitely permeable plate. This is a nontrivial theoretical construction since the normal Casimir force between the plates is now repulsive [10], which can also be regarded as an anomaly [10, 11], due to the technical detail that the discrete wavenumbers in the overlapping space satisfy [10]
In this circumstance the radiation pressure on the left boundary should be
where
It is easy to cast the pressure as
where is a sufficiently strong cutoff function, and its derivatives
After elementary calculations (see Appendix A), we obtain a positive radiation pressure as
One may find, with the information provided by Ref. [10], that this radiation pressure would be in agreement with the negative pressure occurring in this misaligned system. It reveals that the ordinary radiation in the quantum vacuum could also generate the negative pressure required by a constant vacuum energy density.
Before we end this study, a technical note may be given here. In our discussions, the plane wave is assumed to be reflected by the single cross section at the edge, and then possible disturbance to the propagation of the incident field near the overlapping edge is not considered. Is this simplified treatment well-defined? The answer is affirmative. There are two ways to understand this conclusion: 1) From the conventional viewpoint of zero-point energy, Ref. [5] pointed out that the tangential force experienced by the plates in Fig. 1 is irrelevant to the details of the quantum vacuum in the edge regions. This fact suggests that it will not be necessary to complicate our discussion on this point; 2) But a concise proof may also be outlined as follows. Notice that one can always assume an arbitrary curved surface outside (or inside) the space between the plates, which can form a closed surface with the cross section at the edge (including formally the “ends” of the infinite edge along -direction, which is actually an insignificant detail), and near this imaginary surface the propagation of the virtual photons is not disturbed by the edge. Each element of the surface should experience a net, normal momentum flux of the incident field. Then, the incoming flux of momentum through the curved surface is equal to the undisturbed flux through the edge cross section, and it would be simple to realize that this fact should eventually approve our treatment.
Appendix A
We present here a calculation of Eq. (17). The equation may be rewritten as
where
By the substitution and integration by parts,
The third term in the second line in Eq. (26) is an arbitrary constant, which will be determined by a particular choice of the cutoff function on the entire domain and thus is of no physical interest to us. We may get rid of this constant by adding a counterterm to the integral. After the renormalization,
It may be worth noting in this appendix that, from the expressions of virtual photon radiation pressures in the misaligned system under consideration, a desired relation between the normal pressure (see Ref. [7]) and the tangential pressure can be directly obtained as
using a substitution
for each mode, where is the zero-point energy density difference between the continuous and the discrete wavenumbers.
[9]
The zero-point momentum occurs in the quantization of the momentum operator of a field, which, like the zero-point energy, has its roots in the equal-time commutation relations of field operators, see, e.g., C. Itzykson and J.-B. Zuber, Quantum field theory, New York: McGraw-Hill, 1980, p115.